# Retrieve, Merge, Predict: Augmenting Tables with Data Lakes

### v6

@@ Appendix A Construction of YADL
We construct YADL by first reshaping the RDF triplets into binary tables. Then, we group subjects by their Wordnet type and construct 1015 wide-form relational “seed tables” by joining binary tables on the column that contains the triplet subjects.
Each seed table is then partitioned into “sub-tables” by selecting subsets of columns and projecting over them. Different partitioning strategies allow us to vary the characteristics of the resulting sub-tables (i.e., the number of rows and columns of each table), and the size of the resulting data lake.
@@
Transforming triplet tables into wide format tables leads to tables with null values because not all entities of a given type have the same relations. For example, column “hasPopDensity” contains null values for subject “Paris” in  Table 4. The fraction of missing values in the transformed tables is reported in Table 1.
@@ B.2 Selection
The reference selector is No Join, where the ML model is trained directly on the training split of the base table.
Highest Containment Join ranks candidates by exact containment, after measuring it for each candidate. The top-$1$ candidate is joined; ties are broken by taking one candidate at random.
Best Single Join iterates over each candidate one at a time, performs the aggregation and join, then trains an ML model on the joined table. The candidate with the highest model performance on the validation split is selected and joined.
Top-$k$ Full Join aggregates and joins the first $k$ candidates as ranked by a retrieval method, without performing any re-ranking.
Full Join aggregates and joins all candidates from the pool.
Stepwise Greedy Join aims to add more information than what is available in a single table, while reducing the amount of noise added by joining irrelevant tables. The baseline performance is measured by training a model on the base table’s training split and validating on the validation split; candidates are re-ranked by their containment before iterating over each of them. During each iteration, the candidate is aggregated and merged separately with each of the two splits of the current table; a new model is trained on the training split and validated on the validation split. If the performance improves compared to the previous iteration, the newly joined table becomes the new current table, otherwise the candidate is discarded.
@@ B.3 Aggregation
Figure 7: Example of how a left join would duplicate rows from the base table.
@@ C.1 Choosing which baselines to use
Including recent publications often turned out to be very difficult, as the code was either unavailable, or not suited to extensive experiments on a data lake of the size that we investigate in our work. We worked on including Starmie (Fan et al., 2023b), and (partially) Lazo (Fernandez et al., 2019). We were able to test Starmie on some of the data lakes (Binary, YADL Base, YADL 10k) though we did have to adapt and optimize certain sections of the discovery code that did not scale well to our environment. We were unable to run Starmie on the larger YADL 50k, and the less structured Open Data US. Starmie results are illustrated in Section 5. The modified Starmie repository is available here: https://github.com/rcap107/starmie, and we are contributing it upstream. These modifications are code optimizations and fixes for corner cases; they do not change the logic of Starmie.
@@ C.2 Experimental setup
Retrieval We use a containment threshold of 0.2 for the preparation of the MinHash index, and clamp the number of candidates returned by each retrieval method to 30. These values were chosen to balance execution time and expected number of candidates given the distribution of containment encountered in the different data lakes (Figure 8). For Starmie, we use the default parameters defined in the original repository.
@@ C.4 Aggregated results
Figure 12: Pareto frontier analysis for the pipeline steps.
@@
Stepwise Greedy Join is much slower compared to all other methods. This is not surprising, since this selector executes a join and trains a model in each iteration, then re-trains the model at the end of the fit step. In practice, while this added complexity brings an improvement of about 1.6% with respect to Best Single Join, this comes at the cost of a 2.9x increase in run time; even worse, this does not achieve the best performance as Full Join is both faster and slightly better in prediction performance. Highest Containment is significantly faster than all other solutions because it only re-ranks candidates before joining the top-$1$ candidate.
@@
An important observation is that all selectors rely on the candidates proposed by a retrieval method: if these candidates have poor quality, the selectors cannot compensate for that.
@@
Aggregation experiments do not include Full Join and Stepwise Greedy Join because DFS ran out of resources with those selectors. In fact, while DFS brings some benefit in its generation of new features (up to 2.46% in median wrt Any), it is also extremely slow (up to 6x wrt the simpler method), and this was in a simpler case where only one table was joined at a time. With multi-table selectors, this problem became even more noticeable and prevented us from testing the multi-table selectors. This very large memory cost is consistent with Cvetkov-Iliev et al. (2023).
@@
The performance difference between CatBoost and all other methods already highlighted in Figure 2 and Figure 11 is confirmed in Figure 15(d), which shows CatBoost outperforming the second best model in RealMLP by about 21.2%, and the worse NN model in RidgeCV by 26%. Interestingly, CatBoost is also faster than all other methods by a significant amount, likely owing to the optimized preprocessing of categorical variables. RealMLP is the better performing NN model, and has shown a good resilience in the very challenging environment that we are considering; however, it is much slower than all other methods, even when run on GPUs, requiring up to 32x as much time as CatBoost.
@@ C.5 Smart aggregation brings some benefit at a major cost
Figure 13: Aggregation trade-offs relative to run time.
Figure 14: Aggregation trade-offs relative to peak RAM.
@@
Figure Figure 13 and Figure 14 confirm that, while DFS improves the prediction performance, its overall cost remains problematic.
Figure 15: Experimental results across all data lakes, comparing the performance difference between evaluated methods in different steps of the pipeline. The median difference is reported on the right of each plot. (a) compares join selectors, (b) aggregation methods, and (c) ML models. In all cases, results are relative to the “average method”. Starmie results are not reported in this plot as we could not run it on the larger data lakes.
@@ C.7 Execution time breakdown
Figure 18: Relative prediction performance (median on the right) obtained selecting the (top-$1$) candidate according to the ranking provided by the retrieval methods. YADL 50k and Open Data are not included.
@@
C.8 Effectiveness of top-$1$ join with different retrieval methods
Figure 18 reports the prediction performance across different data lakes obtained by joining each base table on the top-$1$ candidate according to each retrieval method’s internal ranking (here, Exact is equivalent to top-$1$ Highest Containment Join). Unsurprisingly, MinHash is far worse than all other methods (up to 13.5% in median) due to the lack of an inherent ordering, and budget constraints; Hybrid MinHash successfully mitigates some of the issues of MinHash, however the imperfect retrieval performance penalizes it. It appears that the first candidate retrieved by Starmie is not as good as what is selected by Exact Matching. We observe that the performance of Starmie drops on YADL 10k: this is likely due to the larger width of the candidate tables, which affects the cosine similarity, and thus worsens the ranking. However, the situation improves for multi-table selectors, suggesting that this issue is mitigated by selecting multiple candidates at once.
@@ C.8 Effectiveness of top-$1$ join with different retrieval methods
C.9 Trade-off
@@
Figure 5 is prepared assuming that query retrieval time increases linearly with the size of a table and that the cost of creating the MinHash index is fixed for the data lake at hand; we use Table 10 as reference to build the figure. We observe that the MinHash indexing cost pays itself off after as few as three queries on average thanks to the fast query time; Hybrid MinHash requires more queries to break even due to its slower query time, yet it remains faster than Exact Matching when the number of query columns is larger. While the assumptions may not hold in general, the plot gives a reasonable estimate of the break-even points.
Starmie scales much worse than other methods as it has a much slower build and query time, and a much larger memory footprint (Figure 5, Table 6).

### v7

@@ Appendix A Construction of YADL
We construct YADL by first reshaping the RDF triplets found in YAGO into binary tables, then, depending on the construction strategy, we assemble new YADL variants by joining binary tables on each other.
@@
Transforming triplet tables into wide format tables introduces null values because not all entities of a given type have the same relations. For example, column “hasPopDensity ” contains null values for subject “Paris” in  Table 4. The fraction of missing values in the transformed tables is reported in Table 1.
@@ B.2 Candidate selection
Highest Containment Join
We first consider a simple method that assumes that no ranking is provided in the previous step. Highest Containment Join ranks candidates by exact containment, after measuring it for each candidate. The top-$1$ candidate is joined; ties are broken by taking one candidate at random. This selector re-ranks candidates by containment even if they have already been ranked by the retrieval method, thus it effectively ignores the previous ranking. Due to its simplicity, Highest Containment Join is very quick to execute, but it is likely to select suboptimal candidates.
Best Single Join
Best Single Join iterates over each candidate one at a time. At each step, it joins the base table on the current candidate, performs the aggregation, trains a ML model, then evaluates the model on a held-out validation step. After all candidates have been evaluated, the best candidate and associated ML model are used: the ML model is re-trained over the entire training split and selected for the final evaluation. The Best Single Join is an exhaustive and expensive method.
Full Join
The Full Join selector does not include any logic for selecting candidates. Given the base table, the Full Join first aggregates candidates, then joins all of them on the base table. Depending on the budget that is provided and the number of columns in the candidates, this may increase the final number of features in the augmented table by a lot. In most of our experiments, we fix the number of candidates to 30. After joining, the ML model is trained over the fully augmented table. Since training is performed only once, this method is faster than the exhaustive selection methods (Best Single Join and Stepwise Greedy Join), but the large number of resulting features may increase the memory pressure.
Stepwise Greedy Join
Stepwise Greedy Join is an extension of Best Single Join that aims to add more information than what is available in a single table, while reducing the amount of noise added by joining irrelevant tables. Firstly, the baseline performance is measured by training a model on the base table’s training split and validating on the validation split. The base table then becomes the “current table” at the first step of algorithm. Then, candidate selection is performed. Candidates are initially re-ranked by their containment, then each candidate is evaluated sequentially. In each iteration, a candidate is joined on the current table, evaluated on the validation split, then depending on whether it improves performance or not it may be kept or discarded. Any time a candidate is retained, its features will be added to the “current table”. Candidates that do not improve performance are discarded. The training procedure of Stepwise Greedy Join is expensive, as the training operation is repeated multiple times, and slows down as the number of features increases. It has however the advantage of avoiding the addition of suboptimal features. Algorithm 1 reports the steps required.
Algorithm 1 Pseudocode of the Stepwise Greedy Join selector.
0:  base table $T$, candidate set $\mathcal{C}~{}=~{}\{c_{1},c_{2}\ldots c_{n}\}$, scoring function $f$
0:  chosen candidates set $\mathcal{S}~{}=~{}\emptyset$
0:  predictor ML
  $T_{train}$, $T_{valid}$ = get_training_splits($T$)
  ML.fit($T_{train}$)
  score=$f$(ML.predict($T_{valid}$))
  $t^{c}_{train}$ = $T_{train}$
  $t^{c}_{valid}$ = $T_{valid}$
  $\mathcal{C}$ = rerank($\mathcal{C}$)
  for all $i$ in $n$ do
     aggregate($c_{i}$)
     $t^{i}_{train}$ = $t^{c}_{train}$ LEFT JOIN $c_{i}$
     $t^{i}_{valid}$ = $t^{c}_{valid}$ LEFT JOIN $c_{i}$
     ML.fit($t^{i}_{train}$)
     $score_{i}$=$f$(ML.predict($t^{i}_{valid}$))
     if $score_{i}$ > $score_{c}$ then
        $score_{c}$ = $score_{i}$
        $t^{c}_{train}$ = $t^{i}_{train}$
        $t^{c}_{valid}$ = $t^{i}_{valid}$
        $\mathcal{S}~{}=~{}\mathcal{S}~{}\cup~{}\{c_{i}\}$
     else
        discard($c_{i}$)
  $t$ = $T$
  for all $s_{i}~{}\in~{}\mathcal{S}$ do
     $t$ = $t$ LEFT JOIN $s_{i}$
  ML.fit($t$)
Single-table selectors and multi-table selectors
@@ B.3 Aggregation
Figure 6: Example of how a left join would duplicate rows from the base table, and how aggregation keeps the same number of samples as the base table (top left).
@@ C.1 Choosing which baselines to use
Including recent publications often turned out to be very difficult, as the code was either unavailable, or not suited to extensive experiments on a data lake of the size that we investigate in our work. We worked on including Starmie (Fan et al., 2023b), and (partially) Lazo (Fernandez et al., 2019). We were able to test Starmie on some of the data lakes (Binary, YADL Base, YADL 10k) though we had to adapt and optimize certain sections of the discovery code that did not scale well to our environment. We were unable to run Starmie on the larger YADL 50k, and the less structured Open Data US. Starmie results are illustrated in Section 5. The modified Starmie repository is available here: https://github.com/rcap107/starmie, and the improvement was merged in the main repository. These modifications are code optimizations and fixes for corner cases; they do not change the logic of Starmie.
@@ C.2 Experimental setup
Figure 8: Schema of the cross-validation setup used in the training pipeline.
Figure 9: Pre-processing steps applied to input tables before being fed to the parametric ML algorithms. Catboost does not require preprocessing. The names reported here reflect the corresponding scikit-learn (Pedregosa et al., 2011) objects.
@@
Retrieval We use a containment threshold of 0.2 for the preparation of the MinHash index, and clamp the number of candidates returned by each retrieval method to 30 (except when specified otherwise). These values were chosen to balance execution time and expected number of candidates given the distribution of containment encountered in the different data lakes (Figure 7). For Starmie, we use the default parameters defined in the original repository.
@@ C.4 Aggregated results
Figure 13: Pareto frontier analysis for the pipeline steps on all data lakes. The prediction performance ($y$-axis) is plotted against retrieval + run time (top) and peak RAM usage (bottom). Time for offline retrieval preparation for MinHash is not reported here. This figure includes all data lakes, but no results with Starmie.
@@
Figure 14: Experimental results across all data lakes, comparing the performance difference between evaluated methods in different steps of the pipeline. The median difference is reported on the right of each plot. (a) compares join selectors, (b) aggregation methods, and (c) ML models. In all cases, results are relative to the “average method”.
@@
Stepwise Greedy Join is much slower compared to all other methods. This is not surprising, since this selector executes a join and trains a model in each iteration, then re-trains the model at the end of the fit step. In practice, while this added complexity brings an improvement of about 1.6% with respect to Best Single Join, this comes at the cost of a 2.1x increase in run time; even worse, this is not enough to beat Full Join in all cases, as the latter method is both faster and slightly better in prediction performance. Highest Containment is significantly faster than all other solutions because it only re-ranks candidates before joining the top-$1$ candidate.
@@
It is important to mention that the picture is more complex than what can be observed exclusively from Figure 14. In fact, Figures 19 and 20 show that, while Full Join appears to be better on average, there are configurations of hyperparameters in which it performs worse than Stepwise Greedy Join due to the addition of unrelated features.
A final important observation is that all selectors rely on the candidates proposed by a retrieval method: if these candidates have poor quality, the selectors cannot compensate for that.
@@
Aggregation experiments do not include Full Join and Stepwise Greedy Join because DFS ran out of resources with those selectors. In fact, while DFS brings some benefit in its generation of new features (up to 2.46% in median wrt Any), it is also extremely slow (up to 4.42x wrt the simpler method), and this was in a simpler case where only one table was joined at a time. With multi-table selectors, this problem became even more noticeable and prevented us from testing the multi-table selectors. This very large memory cost is consistent with Cvetkov-Iliev et al. (2023).
@@
The performance difference between CatBoost and all other methods already highlighted in Figure 2 and Figure 12 is confirmed in Figure 14(d), which shows CatBoost outperforming the second best model in RealMLP by about 21.2%, and the worse NN model in RidgeCV by 26%. Interestingly, CatBoost is also faster than all other methods by a significant amount, likely owing to the optimized preprocessing of categorical variables. RealMLP is the better performing NN model, and has shown a good resilience in the very challenging environment that we are considering; however, it is much slower than all other methods, even when run on GPUs, requiring up to 17x as much time as CatBoost.
@@ C.5 Smart aggregation brings some benefit at a major cost
Figure 15: Pareto frontier analysis testing different aggregation methods. The prediction performance ($y$-axis) is plotted against retrieval + run time (top) and peak RAM usage (bottom). Each row presents the same results, broken down by retrieval method (left), join selector method (center) and predictor (right). Starmie, Full Join and Stepwise Greedy Join are not reported here.
@@
Figure 15 confirms that, while DFS improves the prediction performance, its overall cost remains problematic.
@@ C.7 Execution time breakdown
C.8 Computational trade-offs
Figure 18 is prepared assuming that query retrieval time increases linearly with the size of a table and that the cost of creating the MinHash index is fixed for the data lake at hand; we use Table 10 as reference to build the figure. We observe that the MinHash indexing cost pays itself off after as few as three queries on average thanks to the fast query time; Hybrid MinHash requires more queries to break even due to its slower query time, yet it remains faster than Exact Matching when the number of query columns is larger. While the assumptions may not hold in general, the plot gives a reasonable estimate of the break-even points. S͡tarmie scales much worse than other methods as it has a much slower build and query time, and a much larger memory footprint (Figure 22, Table 6).
@@ C.8 Computational trade-offs
C.9 Interplay between different variables
In Figures 20 and 19 we report the interplay between each variable (Predictor, Retrieval method, Selector, Data lake, Base table) with each other. Each row in a figure is a boxplot that represents a specific variable, while each column on the row breaks down the results by a different variable. We provide these figures to showcase how specific configurations may have behaviors that are hidden by the necessity to aggregate information for the main body. Some interesting takeaways that emerge from this representation are the following:
  • RealMLP Holzmüller et al. (2024) is more resilient than all other parametric methods even in the challenging configurations we are considering.
  • RidgeCV is even more penalized by larger tables (see the results with Full Join and on Open Data)
  • Full Join and Stepwise Greedy Join have similar performance, but Full Join has worse worst-case performance on some data lakes. This confirms the fact that when noise is more prevalent, Full Joining is not a good strategy.
  • Overall, Starmie does not bring a lot of benefits and behaves similarly to Exact Matching in most configurations.
  • Other than Catboost for predictors, no other method appears to clearly dominate its category.
Figure 19: Effect of different variables on the (absolute) prediction performance. Each row reports one variable (e.g., Predictor) and breaks it down by another variable on each column (e.g., by Retrieval method). This plot does not include YADL50k and Open Data US.
Figure 20: Effect of different variables on the (absolute) prediction performance. Each row reports one variable (e.g., Predictor) and breaks it down by another variable on each column (e.g., by Retrieval method). This plot does not include Starmie.
C.10 Effect of top-k
In this section we give provide additional context on the impact of $k$ on the pipeline. Figure 22 shows that the prediction performance saturates quickly as the number of candidates increases, while the memory footprint increases sharply: this behavior is similar to the impact on runtime shown in Figure 4.
Figures 25, 25, and 25 show how increasing the value of $k$ affects the prediction performance, fold runtime, and peak RAM usage respectively as a function of the data lake and base table. We use the reference configuration to obtain these results. The data lake has a large impact on the pipeline, with Open Data and YADL10k having a much larger RAM footprint. This may partially be an artefact of how YADL10k is generated; in Open Data’s case, this is likely due to the fact that some of the tables in the data lake contain potentially hundreds of columns.
Figure 21: Pareto plot showing how the performance changes as a function of the peak fit RAM for different values of $k$.
Figure 23: Prediction performance as a function of the number of candidates obtained in the retrieval step, broken by data lake.
Figure 24: Fold runtime in seconds as a function of the number of candidates obtained in the retrieval step, broken by data lake.
Figure 25: Peak RAM required at fit time as a function of the number of candidates obtained in the retrieval step, broken by data lake.